# Understood: Real-Time Communication Support for Adults with ADHD Using Mixed Reality

### v1

@@ 4.2.2. Real-time Conversation Support
This design serves two main purposes. First, to minimize distractions during conversations, support panels are not constantly visible – they appear only when needed, helping users stay focused on the interaction. Moreover, when multiple panels are active, gaze-based dimming ensures users can concentrate on one feature without overwhelming their field of view (F8, DG5). Second, in face-to-face communication, maintaining appropriate eye contact is crucial for social comfort. By positioning support panels directly in front of the user, the system allows users to seek help discreetly, preserving natural conversational flow while providing meaningful assistance (F3, DG5).
@@
The Summarization Panel includes two components: self-summarization (Figure 5-6) and other-summarization (Figure 5-7), which provide concise recaps of the user’s own speech and their conversational partner’s speech, respectively. Considering the immediacy of dialogue, the system automatically updates these panels after each utterance. At each update, it first extracts the most recent user utterance and then appends the last available summary from the conversation history. This combined input is submitted to GPT-4o (gpt4o) for subsequent processing (the prompt can be seen at subsection C.1). The generated output consists of 4 to 12 keyword-style summary terms displayed in the corresponding panel. These keywords act as memory cues, helping users track the conversational flow and recall key points, especially useful during lapses in attention or external distractions (F4, DG2).
@@ 4.2.3. Status Feedback
Figure 8. User study and procedure for evaluating the effectiveness and usability of Understood. For clarity, the procedures for conversation task (with and without Understood) are presented separately at the bottom. The study involved both patients and clinicians, where patients completed a tutorial, two conversation tasks, questionnaires, and interviews, while clinicians reviewed video recordings and provided evaluations.
@@ 5.2. Procedure
Informed by prior research (EVENSIMKIN202466; barkley2007adhd-ch13; kofler2019working; zulfikar2024memoro; wu2024trinity) and developed in collaboration with clinicians C1 and C2, we designed the user study procedure illustrated in Figure 8. We began by obtaining informed consent and collecting basic demographic information, with a focus on participants’ communication-related challenges. Each participant was then introduced to the Understood system and received a guided walkthrough of its key functionalities while wearing the HoloLens device. Following the tutorial, participants were given 10 minutes to freely explore the system and engage in informal conversation with the researchers, providing them an opportunity to familiarize themselves with the interface and features in a relaxed setting.
@@ 5.2.2. Post-Task Measures and Clinician Interview
Upon completing both conversation tasks, participants were asked to fill out a post-task questionnaire evaluating their experience in terms of efficiency and usability. Responses were recorded using a 5-point Likert scale. The questionnaire items were adapted from well-established instruments, including the System Usability Scale (SUS) (brooke1996sus), the NASA Task Load Index (NASA-TLX) (hart1986nasa), and relevant recent studies (zulfikar2024memoro; wu2024trinity; cai2023paraglassmenu). All the question statements in the post-task questionnaire are listed at subsection D.2. In addition, semi-structured interviews were conducted to gather qualitative feedback and suggestions. Each participant received a compensation of \$15 USD for their participation.
@@ 5.3. Results and Analysis
Based on the results presented in Figure 9 and Table 2, along with insights from semi-structured interviews with both patients and clinicians, we derived the following conclusions.
@@ 5.3.1. Effectiveness
In terms of Response Length, Understood generates concise and intuitive prompts (Summarization: $8.14\pm 1.73$ words; Suggestion: $3.52\pm 0.88$ words), which were deemed appropriate by users, as reflected in their perception ratings (Summarization: $4.30\pm 0.78$; Suggestion: $4.70\pm 0.46$). As P5 observed, “The summarization is brief, never exceeding three lines, and the suggestions are concise, requiring only a few words. I can quickly understand the content and use it to steer my next response without spending too much time reading.”
In terms of Response Time, Understood achieved average response times of $1.25\pm 0.40$ seconds for Summarization and $1.95\pm 0.71$ seconds for Suggestion. The longer duration for Suggestion is attributed to the reliance on the full conversation history, which increases the processing load as the conversation progresses, as detailed in subsubsection 4.2.2. In contrast, Summarization builds upon previous summaries, keeping the processing time relatively constant. Overall, participants responded positively to the system’s responsiveness (ratings: Summarization: $4.60\pm 0.66$; Suggestion: $4.10\pm 0.83$). As P3 commented, “Understood responds in a reasonable time frame. Anyone familiar with LLMs would find the timing quite acceptable.”
Regarding Correctness, responses generated by Understood were generally perceived as helpful by most participants (Summarization: $4.10\pm 0.70$; Suggestion: $4.30\pm 1.00$). As P1 remarked, “When my mind went blank, Understood often provided a reasonable hint to help me continue. The Summarization helped me organize the context, while the Suggestion provided a new opportunity.” Several participants also reported that the the responses encouraged them to expand the scope of the conversation. For example, P4 shared, “I was asked to talk about my favorite place, but I couldn’t immediately think of anything. Then the system suggested ‘City Park’, which reminded me of a wonderful picnic experience from a few years ago. Without the suggestion, I wouldn’t have recalled the memory.”
@@
Regarding User Performance, the results suggest that participants benefited from the support of Understood. While no statistically significant differences were observed in the Number of Pauses (With: $15.60\pm 5.28$; Without: $18.90\pm 7.37$) and Number of Off-Topics (With: $5.10\pm 1.45$; Without: $4.80\pm 1.47$), participants showed significant improvements in their ability to recover from these events. Specifically, there was a notable reduction in Average Pause Duration (With: $4.22\pm 0.86$ seconds; Without: $6.17\pm 0.76$ seconds; $p=0.002$) and Average Off-Topic Duration (With: $34.27\pm 9.25$ seconds; Without: $49.12\pm 6.40$ seconds; $p=0.014$). These findings suggest that while Understood may not significantly alter the frequency of pauses or off-topic responses, it effectively facilitates quicker recovery when such events occur.
@@
Clinicians, however, raised concerns about Understood’s focus on symptom management rather than addressing the underlying causes of ADHD-related communication challenges. As C1 noted, “While I appreciate Understood’s ability to help patients manage communication difficulties, I worry that it only addresses the ‘surface-level’ symptoms of ADHD. A symptom-focused approach may have limited long-term impact. Nonetheless, I see potential in how the system could help patients transfer skills from controlled environments to real-world interactions.” This concern is further discussed in subsection 6.2.
@@
Participants responded positively to Understood in terms of enhancing their Concentration on Conversation ($4.50\pm 0.81$), Eye Contact ($4.00\pm 1.00$), and overall Naturalness ($4.40\pm 0.49$). These results indicate that the system effectively supported participants in maintaining conversational focus and engaging more naturally. P9 commented, “With the summarization feature, I no longer worry about forgetting the previous parts of the conversation, which helps me feel more confident and stay focused while talking.” Similarly, P4 noted, “I can seamlessly integrate the summarization and suggestions into my speech, and the words naturally flow out of my mouth.” P7 added, “When my gaze landed on the red-colored trigger, it subtly reminded me that I had gone off-topic, helping me steer the conversation back on track smoothly.”
@@
However, some participants expressed concerns about maintaining eye contact, noting that their gaze often shifted back and forth between the panels. Despite this, clinicians viewed eye contact more positively from an audience perspective. C2 remarked, “From my perspective, I seldom notice their gaze drifting away from the conversational partner, so the effect is quite acceptable.”
@@
Overall, most participants found Understood to be effective in helping them overcome communication challenges, as reflected in the Helpfulness rating of $4.50\pm 0.67$.
@@ 5.3.2. Usability
In terms of Ease of Use ($4.60\pm 0.49$) and Learnability ($4.90\pm 0.30$), all participants agreed that Understood provided an intuitive interaction mechanism that helped them overcome communication challenges. The minimalistic design further facilitated quick mastery of the system. P4 commented, “A clever solution! Understood tackles challenges with simple, well-defined functions. I was able to grasp its features and usage in a short time, which will be especially helpful for first-time users.” This feedback is supported by the Function Integration rating of $4.50\pm 0.50$. Additionally, participants found the Function Selection Phase useful, as it allowed them to customize the system’s functions and interface to better meet their individual needs.
Regarding User Control ($4.00\pm 0.77$), most participants felt that Understood performed as expected. However, some noted that hallucinations by the LLMs could lead to unexpected responses, causing confusion. P6 added, “When I encounter unexpected responses, I have to put more mental effort into processing them, as I need to filter out these responses from my original thoughts.” As a result, although Understood received a high rating for Low Interruption ($4.60\pm 0.49$) due to its minimal and well-defined design, the rating for Low Distraction was lower than other statements ($4.20\pm 0.87$), primarily due to inaccuracies.
Understood also gained consensus among participants for its low mental ($4.60\pm 0.49$) and physical ($4.60\pm 0.66$) demands. P3 remarked, “The summarization and suggestions were incredibly helpful. I no longer need to worry about my personal deficits, as Understood provides support. I can finally relax.” P10 mentioned that triggering auxiliary functions by gaze was physically relaxing: “If you’ve ever interacted with an AR/VR system, you’ll understand how tiring it can be to raise your arms and hands in midair.”
Satisfaction was also highly rated by participants, with a score of $4.60\pm 0.66$. They found the final Feedback Phase both helpful and uplifting. P5 noted, “As the panel displayed the number of times I sought assistance during the conversation, I could use it to evaluate my performance and track my progress. Seeing the number decrease was quite a rewarding experience.” P7 added, “The confetti animation is really fun! It provides further relaxation after a stressful interpersonal conversation.” Overall, most participants expressed their willingness for Frequent Use of Understood ($4.40\pm 0.66$).
@@ 6. Discussion
The concept of Cognitive Offloading describes the process of reducing cognitive load by utilizing external tools or systems (risko2016cognitive; scaife1996external; armitage2020developmental). Understood aligns with this concept by providing real-time support mechanisms—such as summarization, suggestions, and off-topic detection—that help alleviate cognitive strain for users, particularly those with ADHD. These features allow users to focus on the conversation itself while offloading some of the cognitive demands. While this offloading helps users stay engaged and confident in the moment, excessive reliance on such features could hinder cognitive growth and self-regulation (grinschgl2021consequences). Thus, it is crucial to strike a balance between providing cognitive support and fostering cognitive development. If support becomes too frequent or easily accessible, users may skip the cognitive effort required to engage deeply in conversations. Therefore, the system should provide temporary support that gradually fades as users become more adept at independent communication.

### v2

@@ 4.2.1. Auxiliary Functions Selection
[…]
Illustration of the Conversation Phase interface for Understood, split into two sections labeled A and B on a gray background. Section A shows all three panels “[Self-Summarization]”, “[Other-Summarization]” and “[Word Suggestions]” when no gaze directed, indicated by an eye icon away from the panels. Section B mirrors the same summaries but shows a gaze directed at the “[Word Suggestions]” panel, now only highlighting it. The other two un-gazed panels are faded.
@@ 4.2.2. Real-time Conversation Support
Design trade-offs were carefully considered during system development. While placing auxiliary panels in the central field of view may partially obstruct eye contact, this choice was made after exploring and discarding several alternatives during the design workshop(subsection 3.2). One proposed approach involved anchoring auxiliary panels beside the interlocutor using body tracking, but implementing real-time body tracking on HoloLens introduced significant system overhead and would likely cause noticeable latency, affecting usability and study validity. Another alternative involved placing panels in peripheral areas of the visual field, but this led to exaggerated gaze shifts during use, which disrupted natural eye contact and made the user appear disengaged in conversation – an outcome especially undesirable for socially sensitive users.
This design reflects two interrelated considerations. First, to minimize distractions during conversations, support panels are not constantly visible – they appear only when user needed, helping users stay focused on the interaction. Moreover, when multiple panels are active, gaze-based dimming ensures users can concentrate on one feature without overwhelming their field of view (F8, DG5). Second, in face-to-face communication, maintaining appropriate eye contact is crucial for social comfort. By positioning support panels directly in front of the user, the system allows users to seek help discreetly, preserving natural conversational flow while providing meaningful assistance (F3, DG5).
@@
The Summarization Panel includes two components: self-summarization (Figure 5-6) and other-summarization (Figure 5-7), which provide concise recaps of the user’s own speech and their conversational partner’s speech, respectively. Considering the immediacy of dialogue, the system automatically updates these panels after each utterance. At each update, it first extracts the most recent user utterance and then appends the last available summary from the conversation history. This combined input is submitted to GPT-4o (Hurst et al., 2024) for subsequent processing (the prompt can be seen at subsection C.1). The generated output consists of 4 to 12 keyword-style summary terms displayed in the corresponding panel. These keywords act as memory cues, helping users track the conversational flow and recall key points, especially useful during lapses in attention or external distractions (F4, DG2).
@@ 4.2.3. Status Feedback
Figure 9. User study and procedure for evaluating the effectiveness and usability of Understood. For clarity, the procedures for conversation task (with and without Understood) are presented separately at the bottom. The study involved both individuals with ADHD and clinicians, where individuals with ADHD completed a tutorial, two conversation tasks, questionnaires, and interviews, while clinicians reviewed video recordings and provided evaluations.
[…]
Illustration of the user study and procedure for evaluating Understood’s effectiveness and usability, divided into two main sections. The top section outlines the conversation tasks: ADHD individuals and clinicians start with usage tutorials and trials, followed by random/even assignment to “With Understood” or “Without Understood” groups, depicted with light bulbs. Both groups proceed to post-task questionnaires and interviews. The bottom section details the conversation task process (with/without Understood): starting with topic selection and supplement delivery, followed by divergent response generation and collaborative elaboration, leading to convergent summarization, illustrated with icons and flow arrows. Clinicians’ path includes hands-on trials, video recorded recordings review, and interviews.
@@ 5.2. Procedure
Informed by prior research (Even-Simkin, 2024; Barkley et al., 2007; Kofler et al., 2019; Zulfikar et al., 2024; Wu et al., 2024) and developed in collaboration with clinicians C1 and C2, we designed the user study procedure illustrated in Figure 9. We began by obtaining informed consent and collecting basic demographic information, with a focus on participants’ communication-related challenges. Each participant was then introduced to the Understood system and received a guided walkthrough of its key functionalities while wearing the HoloLens device. Following the tutorial, participants were given 10 minutes to freely explore the system and engage in informal conversation with the researchers, providing them an opportunity to familiarize themselves with the interface and features in a relaxed setting.
@@ 5.2.2. Post-Task Measures and Clinician Interview
Upon completing both conversation tasks, participants were asked to fill out a post-task questionnaire evaluating their experience in terms of efficiency and usability. Responses were recorded using a 5-point Likert scale. The questionnaire items were adapted from well-established instruments, including the System Usability Scale (SUS) (Brooke et al., 1996), the NASA Task Load Index (NASA-TLX) (Hart, 1986), and relevant recent studies (Zulfikar et al., 2024; Wu et al., 2024; Cai et al., 2023). All the question statements in the post-task questionnaire are listed at subsection D.2. In addition, semi-structured interviews were conducted to gather qualitative feedback and suggestions. Each participant received a compensation of \$15 USD for their participation.
@@
[…]
Please read the following ”Result and Analysis” section for the results reported in the image.
@@ 5.3. Results and Analysis
Based on the results presented in Figure 10 and Table 2, along with insights from semi-structured interviews with both individuals with ADHD and clinicians, we derived the following conclusions.
@@ 5.3.1. Effectiveness
In terms of Response Length, Understood generates concise and intuitive prompts (Summarization: 8.14±1.738.14$\pm$ 1.738.14 ± 1.73 words; Suggestion: 3.52±0.883.52$\pm$ 0.883.52 ± 0.88 words), which were deemed appropriate by users, as reflected in their perception ratings (Summarization: 4.30±0.784.30$\pm$ 0.784.30 ± 0.78; Suggestion: 4.70±0.464.70$\pm$ 0.464.70 ± 0.46). As P5 observed, “The summarization is brief, never exceeding three lines, and the suggestions are concise, requiring only a few words. I can quickly understand the content and use it to steer my next response without spending too much time reading.”
In terms of Response Time, Understood achieved average response times of 1.25±0.401.25$\pm$ 0.401.25 ± 0.40 seconds for Summarization and 1.95±0.711.95$\pm$ 0.711.95 ± 0.71 seconds for Suggestion. The longer duration for Suggestion is attributed to the reliance on the full conversation history, which increases the processing load as the conversation progresses, as detailed in subsubsection 4.2.2. In contrast, Summarization builds upon previous summaries, keeping the processing time relatively constant. Overall, participants responded positively to the system’s responsiveness (ratings: Summarization: 4.60±0.664.60$\pm$ 0.664.60 ± 0.66; Suggestion: 4.10±0.834.10$\pm$ 0.834.10 ± 0.83). As P3 commented, “Understood responds in a reasonable time frame. Anyone familiar with LLMs would find the timing quite acceptable.”
Regarding Correctness, responses generated by Understood were generally perceived as helpful by most participants (Summarization: 4.10±0.704.10$\pm$ 0.704.10 ± 0.70; Suggestion: 4.30±1.004.30$\pm$ 1.004.30 ± 1.00). As P1 remarked, “When my mind went blank, Understood often provided a reasonable hint to help me continue. The Summarization helped me organize the context, while the Suggestion provided a new opportunity.” Several participants also reported that the the responses encouraged them to expand the scope of the conversation. For example, P4 shared, “I was asked to talk about my favorite place, but I couldn’t immediately think of anything. Then the system suggested ‘City Park’, which reminded me of a wonderful picnic experience from a few years ago. Without the suggestion, I wouldn’t have recalled the memory.”
@@
Regarding User Performance, the results suggest that participants benefited from the support of Understood. While no statistically significant differences were observed in the Number of Pauses (With: 15.60±5.2815.60$\pm$ 5.2815.60 ± 5.28; Without: 18.90±7.3718.90$\pm$ 7.3718.90 ± 7.37) and Number of Off-Topics (With: 5.10±1.455.10$\pm$ 1.455.10 ± 1.45; Without: 4.80±1.474.80$\pm$ 1.474.80 ± 1.47), participants showed significant improvements in their ability to recover from these events. Specifically, there was a notable reduction in Average Pause Duration (With: 4.22±0.864.22$\pm$ 0.864.22 ± 0.86 seconds; Without: 6.17±0.766.17$\pm$ 0.766.17 ± 0.76 seconds; $p=0.002$) and Average Off-Topic Duration (With: 34.27±9.2534.27$\pm$ 9.2534.27 ± 9.25 seconds; Without: 49.12±6.4049.12$\pm$ 6.4049.12 ± 6.40 seconds; $p=0.014$). These findings suggest that while Understood may not significantly alter the frequency of pauses or off-topic responses, it effectively facilitates quicker recovery when such events occur.
@@
Clinicians, however, raised concerns about Understood’s focus on symptom management rather than addressing the underlying causes of ADHD-related communication challenges. As C1 noted, “While I appreciate Understood’s ability to help ADHD people manage communication difficulties, I worry that it only addresses the ‘surface-level’ symptoms of ADHD. A symptom-focused approach may have limited long-term impact. Nonetheless, I see potential in how the system could help ADHD people transfer skills from controlled environments to real-world interactions.” This concern is further discussed in subsection 6.3.
@@
Participants responded positively to Understood in terms of enhancing their Concentration on Conversation (4.50±0.814.50$\pm$ 0.814.50 ± 0.81), Eye Contact (4.00±1.004.00$\pm$ 1.$0$ 1.00), and overall Naturalness (4.40±0.494.40$\pm$ 0.494.40 ± 0.49). These results indicate that the system effectively supported participants in maintaining conversational focus and engaging more naturally. P9 commented, “With the summarization feature, I no longer worry about forgetting the previous parts of the conversation, which helps me feel more confident and stay focused while talking.” Similarly, P4 noted, “I can seamlessly integrate the summarization and suggestions into my speech, and the words naturally flow out of my mouth.” P7 added, “When my gaze landed on the red-colored trigger, it subtly reminded me that I had gone off-topic, helping me steer the conversation back on track smoothly.”
@@
However, some participants expressed concerns about maintaining eye contact, noting that their gaze often shifted back and forth between the panels. Despite this, clinicians viewed eye contact more positively from an audience perspective. C2 remarked, “From my perspective, I seldom notice their gaze drifting away from the conversational partner, so the effect is quite acceptable.” This gaze-shifting behavior may also be interpreted as a user-initiated coping strategy. By consciously shifting their gaze between the panels, participants appeared to regulate their attention and recover from conversational breakdowns more effectively.
@@
Overall, most participants found Understood to be effective in helping them overcome communication challenges, as reflected in the Helpfulness rating of 4.50±0.674.50$\pm$ 0.674.50 ± 0.67.
@@ 5.3.2. Usability
In terms of Ease of Use (4.60±0.494.60$\pm$ 0.494.60 ± 0.49) and Learnability (4.90±0.304.90$\pm$ 0.304.90 ± 0.30), all participants agreed that Understood provided an intuitive interaction mechanism that helped them overcome communication challenges. The minimalistic design further facilitated quick mastery of the system. P4 commented, “A clever solution! Understood tackles challenges with simple, well-defined functions. I was able to grasp its features and usage in a short time, which will be especially helpful for first-time users.” This feedback is supported by the Function Integration rating of 4.50±0.504.50$\pm$ 0.504.50 ± 0.50. Additionally, participants found the Function Selection Phase useful, as it allowed them to customize the system’s functions and interface to better meet their individual needs.
Regarding User Control (4.00±0.774.00$\pm$ 0.774.00 ± 0.77), most participants felt that Understood performed as expected. However, some noted that hallucinations by the LLMs could lead to unexpected responses, causing confusion. P6 added, “When I encounter unexpected responses, I have to put more mental effort into processing them, as I need to filter out these responses from my original thoughts.” As a result, although Understood received a high rating for Low Interruption (4.60±0.494.60$\pm$ 0.494.60 ± 0.49) due to its minimal and well-defined design, the rating for Low Distraction was lower than other statements (4.20±0.874.20$\pm$ 0.874.20 ± 0.87), primarily due to inaccuracies.
Understood also gained consensus among participants for its low mental (4.60±0.494.60$\pm$ 0.494.60 ± 0.49) and physical (4.60±0.664.60$\pm$ 0.664.60 ± 0.66) demands. P3 remarked, “The summarization and suggestions were incredibly helpful. I no longer need to worry about my personal deficits, as Understood provides support. I can finally relax.” P10 mentioned that triggering auxiliary functions by gaze was physically relaxing: “If you’ve ever interacted with an AR/VR system, you’ll understand how tiring it can be to raise your arms and hands in midair.”
Satisfaction was also highly rated by participants, with a score of 4.60±0.664.60$\pm$ 0.664.60 ± 0.66. They found the final Feedback Phase both helpful and uplifting. P5 noted, “As the panel displayed the number of times I sought assistance during the conversation, I could use it to evaluate my performance and track my progress. Seeing the number decrease was quite a rewarding experience.” P7 added, “The confetti animation is really fun! It provides further relaxation after a stressful interpersonal conversation.” Overall, most participants expressed their willingness for Frequent Use of Understood (4.40±0.664.40$\pm$ 0.664.40 ± 0.66).
@@ 6. Discussion
The concept of Cognitive Offloading describes the process of reducing cognitive load by utilizing external tools or systems (Risko and Gilbert, 2016; Scaife and Rogers, 1996; Armitage et al., 2020). Understood aligns with this concept by providing real-time support mechanisms—such as summarization, suggestions, and off-topic detection—that help alleviate cognitive strain for users, particularly those with ADHD. These features allow users to focus on the conversation itself while offloading some of the cognitive demands. While this offloading helps users stay engaged and confident in the moment, excessive reliance on such features could hinder cognitive growth and self-regulation (Grinschgl et al., 2021). Thus, it is crucial to strike a balance between providing cognitive support and fostering cognitive development. If support becomes too frequent or easily accessible, users may skip the cognitive effort required to engage deeply in conversations. Therefore, the system should provide temporary support that gradually fades as users become more adept at independent communication.
@@
6.2. Safeguarding User’s Autonomy through System Usage
During our user study, concerns were raised about participants’ potential over-reliance on Understood, which could lead to a diminished sense of autonomy. Autonomy refers to the ability to make self-directed decisions free from external coercion or normative pressure (Wertenbroch et al., 2020; Kornmehl, 2025; Fink et al., 2024). Recent work has emphasized autonomy as a key factor influencing user trust and engagement (Vantrepotte et al., 2022; Pagliari et al., 2022). For example, Villa et al. (Villa et al., 2024) found that autonomy can be enhanced by explicitly providing users with control and increasing their confidence during conversations.
Understood offers real-time communication support such as summarization and suggestions, intended to help users express themselves more comfortably. However, we recognize that if users follow these suggestions passively, it may create a subtle power imbalance, leading them to defer to the system rather than reflect on their own communicative intent.
To address this concern, we emphasize that Understood is not designed to prescribe any normative communication style. Instead, its functionality is grounded in customizability and user control. In the Function Selection Phase, users choose the types of support they wish to activate based on their personal needs. In the Conversation Phase, support is provided only through explicit user triggering, while visual cues fade over time to reduce intrusiveness and the risk of passive reliance on the system. Finally, the Feedback Phase allows users to reflect on their experience and opt out at any time, thereby avoiding pressure or normative imposition.
That said, while we acknowledge the autonomy risks inherent in any assistive technology, we believe that Understood safeguards autonomy through intentional, reflexive, and user-centered design.